OpenStudy (jpsmarinho):

The magnitude of the velocity of a ball is 10m/s when it is on the high tall point. What's the mag of the velocity of the ball 1 sec after reaches this point?

OpenStudy (fellowroot):

i dont know what high tall point is?

OpenStudy (lgbasallote):

maybe you can use the formula \[V_2 = V_1 + at\] where V2 is the final velocity V1 is initial velocity a is accleration t is time in this case, initial velocitiy is 10 m/s; acceleration is -9.8 m/s^2 and t is 1 second so substituting those into the formula \[V_2 = 10 + (-9.8)(1)\] \[\implies V_2 = 10 - 9.8\] \[\implies V_2 = 0.2\; m/s\] of course, that is just based on this info given....i could be wrong since the info is a little vague

OpenStudy (fellowroot):

yeah tall point does not ring a bell

OpenStudy (lgbasallote):

it can't me maximum height..

OpenStudy (anonymous):

High tall point, to me anyway, signifies a velocity in the y direction of zero. Therefore the velocity in the x direction, if air resistance is neglected, should remain constant. 1 second later the x velocity should be the same, but the y velocity will have changed by 9.8 m/s. Just add the two velocity vectors to find the total magnitude. I think.